The nucleus of thorium isotope ²³⁴90Th undergoes three consecutive alpha decays . Determine the charge of the resulting isotope
1
Expert's answer
2020-08-10T14:34:04-0400

In every alpha-decay the charge of the nucleus decreases by the charge of alpha-particle, so it decreases by 2. The initial charge of the thorium nucleus is 90, so after three alpha decays the charge will be "90-3\cdot 2 = 90-6 = 84", so we'll obtain a nucleus of polonium. The mass number will decrease by "3\cdot4 = 12."